OpenStudy (anonymous):

radical(.1275/x) = .00303 Solve for x.

OpenStudy (anonymous):

first square both sides to get .1275/x = 0.00303^2 then multiply both sides by x to get .1275 = 0.00303^2*x then divide both sides by 0.00303^2 to solve for x. x = .1275/( 0.00303^2) use a calculator to simplify